Please help me with this question. The dimensions of a rectangle are proportional to 3 and 5, and the perimeter is 192 cm. Find the area of the rectangle.
Answer by rothauserc(4718)

Our dimensions are 3x and 5x, therefore we have the following equation for the perimeter,
2(3x) + 2(5x) = 192
6x + 10x = 192
16x = 192
x = 12
therefore our two dimensions are 3*12 = 36 and 5*12=60
the area of the rectangle is 36 * 60 = 2160 square cm's